a traveler is walking on a moving walkway in an airport. the traveler must walk back on the walkway to get a bag he forgot. the traveler's ground speed is 2 feet per second against the walkway and 6 feet per second whith the walkway. what is the traveler's speed off the walkway? what is the speed of the moving walkway.
if you showed the work it would be very appreciated
thanks :)

Answer by ankor@dixie-net.com(22740)

let t = traveler's normal walking speed
let w = walkway speed
:
Write an equation for each scenario
t - w = 2
t + w = 6
---------------adding eliminates w, find
2t = 8
t = 4 ft/sec, travelers speed
and
4 + w = 6
w = 6 - 4
w = 2 ft/sec, walkway speed